Please use clear handwriting and circle answers thanks.

3. -/1 pointsOSColPhys2016 3.3.WA.014 While in a park, you walk west for 52 m, then you walk 27.0° north of west for 45 m, and finally you walk due north for 25 m. Find the components of your final displacement, from your initial to final point, along the north and west directions. (a) displacement component due north (b) displacement component due west
Answer #1

di = -52 îm d = _ us cos27° î tus sinaro Î - = -40.1 î t 20-4 ĵ m 2 = 25 9 m î tusou î m Now ă = a + a + d₂i =-92.1 (a) (b) 4